• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name b raue

    Log entry time 07:19:14 on April23,2001

    Entry number 61101

    keyword=white scan shift summary


    White run continued. Finishing with 44.3 degree running.

    Run momentum target start stop events comments
    3816 1.0437 LH2 23:47 0:11 1.8M
    3817 1.0228 LH2 0:22 0:47 2.0M
    3818 1.0228 LH2 0:55 1:11 2.0M Open slit running
    3819 1.0228 LH2 1:17 1:40 1.8M Sieve slit running
    3820 1.0228 optics 1:45 2:05 370k optics target
    3821 1.0023 LH2 2:21 2:45 2.0M back to 6 msr slit
    3822 1.0023 LH2 2:49 3:07 1.3M Open slit,
    RF problems, down for 2:42
    3823 1.0023 LH2 5:49 5:56 950k Open slit
    3824 0.9823 LH2 6:07 6:26 2.0M Open slit
    3825 0.9823 LH2 6:29 6:51 2.0M 6 msr slit
    3826 0.9823 optics 6:57 7:17 400k

    PROGRAM COMPLETE!
    Going to controlled access